OpenStudy (anonymous):

what is the derivative of (5x^(2)-4)/2xsqrt(x) ?

OpenStudy (anonymous):

\[(5x^2-4)\div(2x \sqrt{x})dx\]

OpenStudy (anonymous):

ick your math teacher must hate you

OpenStudy (anonymous):

actually if it was me, rather than using the quotient rule i would divide each term by \[2x^{\frac{3}{2}}\] and then use the power rule on \[\frac{5}{2}x^{\frac{1}{2}}-2x^{-\frac{1}{2}}\]should be straight forward

OpenStudy (anonymous):

um, yes. The real question is to use the fundamental theorem of calculus to evaluate the following.. \[\int\limits_{4}^{9}(5x^2-4)\div(2x \sqrt{x})dx\]

OpenStudy (anonymous):

oh then forget what i wrote, you don't want the derivative you want the "anti derivativef"

OpenStudy (anonymous):

\[\int \frac{5}{2}x^{\frac{1}{2}}-2x^{-\frac{3}{2}}dx\]

OpenStudy (anonymous):

now use the power rule backwards to find the anti derivative, plug in the numbers etc

OpenStudy (anonymous):

ohhh, well no wonder i wasn't getting the right answer! haha. i'm so new at this, i jumped into calculus without taking any precalculus at all, stupid move on my behalf.

OpenStudy (anonymous):

can you show me the steps? i have to do 6 other questions very similar to this, i'm sorry.